4|2y-7|+5=9 can you help me find the value of y and show me the steps?
Answer by rothauserc(4718)

4|2y-7|+5=9
4|2y-7| = 4
|2y-7| = 1
CASE 1
2y-7 = 1
2y = 8
y = 4
CASE 2
-(2y-7) = 1
-2y +7 = 1
-2y = -6
y = 3
SOLUTIONS
y = 4 or y = 3